OpenStudy (anonymous):

give medal! 1. Solve using linear combination. y – x = –3 y + x = 3 (Points : 1) (–3, 0) (3, 0) (0, 3) (0, –3)

OpenStudy (anonymous):

y-x=-3 y+x = 3 add them to eliminate the x ----------------- 2y = 0 y = 0 y-x=-3 y+x = 3 subtract them to eliminate y ----------------------- -2x = -6 x = 3 so the coordinates would be (3, 0)